OpenStudy (rashadocmuion):

Explain and Answer Please :) Transform the following Equation written in standard form into slope-intercept form: 2x+3y=5 A. y=(3/2)x+5 B.y=-(2/3)x+5 C.y=-2x+5 D.y=-(2/3)x+5/3

OpenStudy (rashadocmuion):

@faiithhh

OpenStudy (rashadocmuion):

@Preetha

OpenStudy (rashadocmuion):

i think its D

OpenStudy (calculusxy):

you would want to isolate y because in the slope-intercept form you have y = mx + b move 2x to the other side of the equation 3y = 5 - 2x divide both sides by 3 to isolate y y = 5/3 + (-2/3)x or it can be rearranged as `y = -(2/3)x + 5/3`

OpenStudy (calculusxy):

so yes you're correct with the answer choice of D

OpenStudy (rashadocmuion):

Thank you :)